Vicky is 5 years older than Pam. If the sum of their ages is 67 years, how old is each?
Answer by ankor@dixie-net.com(22740)

Write an equation for each statement
:
Vicky is 5 years older than Pam.
v = p + 5
If the sum of their ages is 67 years, how old is each?
v + p = 67
:
Use elimination here, rearrange the 1st equation
v - p = 5
v + p = 67
-----------adding eliminates p, find v
2v = 72
v = 72/2
v = 36 is Vicky's age
:
I"ll let you find Pam's age